OpenStudy (anonymous):

Solve each simultaneous equation X-Y=73 X-2Y=4

OpenStudy (anonymous):

subtract the two

OpenStudy (anonymous):

Can you show me?

OpenStudy (unklerhaukus):

i X-Y = 73 ii X-2Y = 4 i-ii (X-Y)-(X-2Y) = (73)-(4) Y =

OpenStudy (anonymous):

okay ummm this format you put it in is really confusing... but i understand what you did. Sooo whats next?

OpenStudy (unklerhaukus):

did you find Y ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

If you use substitution, you can solve the first equation in terms of one variable and solve for the other. So: x - y = 73 (+y) (+y) x = 73 + y Then you have x - 2y = 4 Since you have a value for x in terms of y, you can also write the equation as (73 + y) - 2y = 4 If you solve for y, you will get a number you can substitute for y in one of the other equations.